Find the minimum number of workstations rounded to the nearest 10ths place. Demand = 1,104 widgets per month Available time: 6 hours in a typical work day. 6 days in a typical work week. 4 weeks in a typical work month.

tasks Time(min) Precedence

1 5

2 4 1

3 2.7 1

4 2.3 2

5 5.7 3

6 3 2,3
Answer #1

Determine total time available:

Total time available = weeks/month x days/week x hours/day x min/hour

Total time available = 4 x 6 x 6 x 60 = 8640 mins/month

Demand = 1104 Units / month

2. Determine cycle time

Cycle time = time available / demand = 8640 / 1104 = 7.8261 mins/unit

Total task time = sum of all task time = 5+4+2.7+2.3+5.7+3 = 22 7 mins

Determine no. Of workstation required:

No. Of work station required = total task time / cycle time = 22.7/7.8261

N = 2.9

Minimum no. of workstation required = 3


answered by: ANURANJAN SARSAM